# Provably Efficient Multi-Objective Bandit Algorithms under Preference-Centric Customization

### v1

@@ 1 Introduction
Multi-objective multi-armed bandit (MO-MAB) problem is an important extension of the multi-armed bandits (MAB) [1]. In MO-MAB problems each arm is associated with a $D$-dimensional reward vector. In this environment, objectives could conflict, leading to arms that are optimal in one dimension, but suboptimal in others. A natural solution is utilizing Pareto ordering to compare arms based on their rewards [1]. Specifically, for any arm $i\in[K]$, if its expected reward $\boldsymbol{\mu}_{i}$ is non-dominated by that of any other arms, arm $i$ is deemed to be Pareto optimal. The set containing all Pareto optimal arms is denoted as Pareto front $\mathcal{O}^{*}$. Formally, $\mathcal{O}^{*}=\{i\mid\boldsymbol{\mu}_{j}\not\succ\boldsymbol{\mu}_{i},\forall j\in[K]\setminus i\}$, where $\boldsymbol{u}\succ\boldsymbol{v}$ holds if and only if $\boldsymbol{u}(d)>\boldsymbol{v}(d),\forall d\in[D]$. The performance is then evaluated by Pareto regret, which measures the cumulative minimum distance between the learner’s obtained rewards and rewards of arms within $\mathcal{O}^{*}$ [1]. However, simply obtaining a solution that has good Pareto regret does not take into account the fact that individual users would like to pick the choice that matches their specific needs. As the example depicted in Fig. 1, given multiple Pareto optimal restaurants, one user may give a higher preference to quality, while another user may give a higher preference to affordibility. This means that user preferences need to be accounted for in the MO-MAB problem set up in order to choose the right solution on the Pareto front $\mathcal{O}^{*}$. This is the focus of this paper.
@@
Numerous MO-MAB studies have been conducted but most of them achieve Pareto optimality via an arm selection policy that is uniform across all users, which we refer to as a global policy. Specifically, one representative line of research focuses on efficiently estimating the entire Pareto front $\mathcal{O}^{*}$, and the action is randomly chosen on the estimated Pareto front [1, 2, 3, 4, 5]. Another line of research transforms the $D$-dimensional reward into a scalar using a scalarization function, which targets a specific Pareto optimal arm solution without the costly estimation of entire Pareto front [1, 6, 7, 8]. These studies construct the scalarization function in a user-agnostic manner, causing the target arm solution to remain the same across different users. However, simply achieving Pareto optimality using a global policy may not yield favorable outcomes, since, as mentioned earlier, users often have diverse preferences across different objectives. Consider the following scenario depicted in Fig. 1(a), where two users with distinct preferences interact with a conversational recommender to find a nearby restaurant for dinner. The upper section lists restaurant options, each associated with multi-dimensional rewards (e.g., price, taste, service), while the lower section shows the dialogues and users’ reward ratings for the recommendations. Clearly, restaurants A, B, and C are Pareto optimal, as none of their rewards are dominated by others. Previous research using a global policy would either randomly recommend a restaurant from A, B, or C, or select one based on a fixed global criterion to achieve Pareto optimality. However, while recommending a restaurant like B might lead to positive feedback from user-1, it is likely to result in a low reward rating from user-2, who prefers an economical meal, since restaurant B is expensive. In contrast, Fig. 1(b) illustrates that when the system accurately captures user preferences (e.g., user-1 prefers a tasty meal, while user-2 prefers a cheap meal), it can select options more likely to receive positive reward ratings from both users. Therefore, we argue that optimizing MO-MAB should be customized based on the user preferences rather than solely aiming for Pareto optimality with a global policy.
@@
To fill this gap, we introduce a formulation of MO-MAB problem, where each user is associated with a $D$-dimensional preference vector, with each element representing the user’s preference for the corresponding objective. Formally, in each round $t$, user incurs a stochastic preference $\boldsymbol{c}_{t}\!\in\!\mathbb{R}^{D}$. The player selects an arm $a_{t}$ and observes a stochastic reward $\boldsymbol{r}_{a_{t},t}\!\in\!\mathbb{R}^{D}$. We define the scalar overall-reward as the inner product of arm reward $\boldsymbol{r}_{a_{t},t}$ and user preference $\boldsymbol{c}_{t}$. The learner’s goal is to maximize the overall-reward accrued over a given time horizon. For performance evaluation, we define the regret metric as the cumulative expected gap related to the overall-reward. We term this problem as Preference-Aware MO-MAB (PAMO-MAB).
@@
While interactive user modeling and customized optimization cross multiple objectives present promising experimental results in some areas including recommendation [9], ranking [10], and more [11], there are no theoretical studies on MO-MAB customization under explicit user preferences. Particularly, two open questions remain: (1) how to develop provably efficient algorithms for customized optimization under different preference structure (e.g., unknown preference, hidden preference)? (2) how does the additional user preferences impact the overall performance?
@@
  • We make the first effort to address the open questions above. Motivated by real applications, we consider the PAMO-MAB problem under two preference structures: unknown preference with feedback, and unknown preference without feedback (hidden preference), with tailored algorithms that are proven to achieve near-optimal regret in each case. The expressions of our results are in an explicit form that capture a clear dependency on preference. To the best of our knowledge, this is the first work that explicitly showcases the fundamental impact of user preference in the regret optimization of MO-MAB problems.
  • For the unknown preference case, we propose an near-optimal algorithm PRUCB-UP incorporating two new designs: Preference estimation and Preference-aware optimization, to enable effective preference capture and optimization under preference-centric customization. For regret analysis, we introduce a novel approach using a tunable parameter $\epsilon$ to decompose the suboptimal actions into two disjoint sets, addressing the joint impact of preference and reward estimation errors on regret.
@@
  • For the hidden preference case, we propose PRUCB-HP, a novel near-optimal algorithm that addresses the unique challenges of hidden PAMO-MAB with two key designs: (1) A weighted least squares-based hidden preference learner, with weights set as the inverse squared $\ell_{2}$-norm of reward observations, to resolve the random mapping issue caused by random preferences, and (2) A dual-exploration policy with novel bonus design to balance the trade-off between local exploration for identifying better reward arms and global exploration for refining preference learning.
2 Related Work
Multi-Objective Multi-Armed Bandits. MO-MAB extends scalar rewards in the standard MAB problem to multi-dimensional vectors. The Pareto-UCB [1] introduced the MO-MAB framework and Pareto regret as a metric, achieving $O(\log T)$ Pareto regret using the UCB technique. Other techniques, including Knowledge Gradient [12] and Thompson Sampling [13], have subsequently been adapted for MO-MAB. Additionally, researchers have extended the contextual setup to MO-MAB [2, 3]. These studies aim to approximate the entire Pareto front $\mathcal{O}^{*}$, and employ a random arm selection policy on the estimated Pareto front for Pareto optimality. Considering computing the full Pareto front is expensive, another line of work propose to converts the multi-dimensional reward into a scalar value through a scalarization function, targeting a specific Pareto optimal solution while not the entire Pareto front. Our work also falls within the realm of this framework. The scalarization function can either be randomly initialized (chosen) [1, 8], or optimized based on a fixed metric, such as the Generalized Gini Index score [6, 7]. Nonetheless, existing studies primarily achieve Pareto optimality through a global policy for arm selection across all users. As discussed in Section 1, merely achieving Pareto optimality with a global policy may not yield favorable outcomes, as users have diverse preferences on different objectives.
Preference-based MO-MAB optimization. Recent studies have explored MO-MAB optimization using lexicographic order [14] to reflect user preferences. In lexicographic order, objectives are prioritized hierarchically, where the first objective takes absolute precedence over the second, and so on. citehuyuk2021multi first introduced lexicographic order to MO-MAB, and citecheng2024hierarchize extended it to mixed Pareto-lexicographic environments. However, lexicographic order may not adequately capture a user’s overall satisfaction in real-world applications, where preferences often involve trade-offs rather than strict prioritization. For example, a user may prefer a \$10 meal with good taste over a \$9.5 meal with poor taste, even though cost is a priority. Our work proposes a more general framework that incorporates a weighted order based on the user’s explicit preference space. Notably, the lexicographic order becomes a special case of our proposed PAMO-MAB framework.
@@ 3 Problem Formulation
We consider MO-MAB with $K$ arms and $D$ objectives. At each round $t\in[T]$, the learner chooses an arm $a_{t}$ to play and observes a stochastic $D$-dimensional reward vector $\boldsymbol{r}_{a_{t},t}\in\mathcal{R}\subseteq\mathbb{R}^{D}$ for action $a_{t}$, which we refer to as reward. For the reward, we make the following standard assumption:
Assumption 3.1 (Bounded stochastic reward).
@@
User preferences. At each round $t$, we consider the user to be associated with a stochastic $D$-dimensional preference vector $\boldsymbol{c}_{t}\in\mathcal{C}\subseteq\mathbb{R}^{D}$, indicating the user preferences across the $D$ objectives. We refer to this vector as preference for short. Specifically, we make the following assumptions:
Assumption 3.2 (Bounded stochastic preference).
For $t\in[T],d\in[D]$, each preference entry $\boldsymbol{c}_{t}(d)$ is independently drawn from a fixed distribution (either known or unknown) with mean $\boldsymbol{\overline{c}}(d)$ and variance $\sigma_{c,d}^{2}$, satisfying $\boldsymbol{c}(d)\geq 0$, $\|\boldsymbol{c}_{t}\|_{1}\leq\delta$, $\sigma_{c,d}^{2}\in[0,\sigma^{2}_{c}]$.
Assumption 3.3 (Independence).
For any $t\!\in\![T]$, $i\!\in\![K]$, $d_{1},d_{2}\!\in\![D]$, $\boldsymbol{r}_{i,t}(d_{1})$, $\boldsymbol{c}_{t}(d_{2})$ are independent.
Assumption 3.3 is common in real applications since $\boldsymbol{c}_{t}$ and $\boldsymbol{r}_{t}$ are inherently determined by independent factors: user characteristics and arm properties. For example, an individual user’s preferences do not influence a restaurant’s location, environment, pricing level, etc., and vice versa.
Preference-aware reward. We define an overall-reward as the inner product of arm’s reward and user’s preference, which models the user reward rating under their preferences. Specifically, we refer to the inner product mapping $\Phi:\mathcal{C}\times\mathcal{R}\rightarrow\mathbb{R}$ as the aggregation function. In each round $t$, the overall-reward $g_{a_{t},t}$ for the chosen arm $a_{t}$ is defined as: $g_{a_{t},t}=\Phi(\boldsymbol{c}_{t},\boldsymbol{r}_{a_{t},t})=\boldsymbol{c}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$.
To evaluate the learner’s performance, we define regret as the cumulative difference between the expected overall-reward by selecting the arm with the highest expected overall-reward at each time $t$ and the expected overall-reward under the learner’s policy:
$R(T)=\sum^{T}_{t=1}\mathbb{E}[g_{a^{*},t}-g_{a_{t},t}]=\sum^{T}_{t=1}\boldsymbol{\overline{c}}^{\top}(\boldsymbol{\mu}_{a^{*}}-\boldsymbol{\mu}_{a_{t}})$ (1)
where $a^{*}\!=\!\operatorname*{arg\,max}_{i\in[K]}\boldsymbol{\overline{c}}^{\top}\boldsymbol{\mu}_{i}$ is the optimal arm. The goal is to minimize the regret $R(T)$. We term this problem as Preference-Aware MO-MAB (PAMO-MAB).
@@ 4 A Lower Bound
In the following, we develop a lower bound (Proposition 1) on the defined regret for PAMO-MAB. Such a lower bound will quantify how difficult it is to control regret without preference-adaptive policies under PAMO-MAB. Firstly, we present a definition characterizing a class of MO-MAB algorithms that are "preference-free".
@@
Let $\boldsymbol{\mathrm{c}}^{t}=\{\boldsymbol{c}_{1},\boldsymbol{c}_{2},...,\boldsymbol{c}_{t}\}\in\mathbb{R}^{D\times t}$ be the preference sequence up to $t$ episode with mean $\boldsymbol{\overline{c}}$. Let $\pi_{t}^{\mathcal{A}}$ be the policy of algorithm $\mathcal{A}$ at time $t$ for selecting arm $a_{t}$. Then $\mathcal{A}$ is defined as preference-free if its policy $\pi_{t}^{\mathcal{A}}$ is independent of $\boldsymbol{\mathrm{c}}^{t}$ and $\boldsymbol{\overline{c}}$, i.e., $\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i\mid\boldsymbol{\mathrm{c}}^{t},\boldsymbol{\overline{c}})=\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i)$ for all arms $i\in[K]$ and all episodes $t\in[T]$.
To our knowledge, most existing algorithms in theoretical MO-MAB studies [1, 6, 8, 15, 16] fall within the class of preference-free algorithms, which employ a global policy for arm selection, while neglecting users’ preferences.
@@
Assume an MO-MAB environment contains multiple objective-conflicting arms, i.e., $|\mathcal{O}^{*}|\geq 2$, where $\mathcal{O}^{*}$ is the Pareto Optimal front. Then, for any preference-free algorithm, there exists a subset of preference such that the regret $R(T)=\Omega(T)$.
@@
Proposition 1 shows that for PAMO-MAB problem with $|\mathcal{O}^{*}|\geq 2$, sub-linear regret is unachievable for preference-free algorithms. This is because, for any arm $i\in\mathcal{O}^{*}$ that is optimal in one preference subset $\mathcal{C}^{+}$, there exists another preference subset $\mathcal{C}^{-}$ where arm $i$ becomes suboptimal, while preference-free algorithms cannot adapt their policies to varying preference across the entire space $\mathcal{C}$. Please see Appendix B for the detailed proof. We therefore ask the following question: Can we design preference-adaptive algorithms that achieve sub-linear regret for PAMO-MAB? The answer is yes. In the following, we analyze PAMO-MAB under two scenarios: preference feedback provided and hidden preference, demonstrating that with preference adaptation, sub-linear regret can indeed be achieved.
5 The Case when the Preference Is Provided
As a warm-up, we begin with a simpler case where the user’s preferences are explicitly provided to the agent either before or after decision making. This setup is prevalent in numerous real-world applications. Many online systems now allow users to express their preferences through interactive techniques, such as conversations and prompt design, either before or after taking action. For example, as shown in Fig. 2, a user can explicitly share her movie preferences with an LLM-based chat system either prior to receiving a recommendation ( in the initial input) or after receiving one ( in the follow-up conversation).
We observe that when the preference $\boldsymbol{c}_{t}$ is given before decision-making, by simply using the preference $\boldsymbol{c}_{t}$ as a weight vector on the reward estimate, the problem collapses into a single-objective MAB. Therefore, we focus on the case where the preference is unknown and only revealed after decision-making. Formally, at each round $t$, the learner first selects an arm $a_{t}$, and then observes the reward $\boldsymbol{r}_{a_{t}}$ and user’s preference $\boldsymbol{c}_{t}$. As we will show later, the known preference case can be handled as a special instance within our developed framework for unknown preference.
As discussed in Section 4, policy adapting to user preference is crucial. To achieve this, we develop Preference-UCB for Unknown Preference (PRUCB-UP, presented in Algorithm 1). At a high level, PRUCB-UP is an extension of the UCB approach [17] with two new components.
Preference estimation. Capturing user preferences is a fundamental step toward preference adaptation. Due to the unknown preference expectation, we leverage the empirical average of historical preference feedback as the preference estimate. For $t\geq 1$, preference estimate is updated as
$\hat{\boldsymbol{c}}_{t+1}=\big{(}(t-1)\boldsymbol{\hat{c}}_{t}+\boldsymbol{c}_{t}\big{)}/t.\vspace{-3pt}$ (2)
$\textstyle N_{i,t+1}$ $=N_{i,t}+\mathds{1}_{\{a_{t}=i\}},$ (3)
$\hat{\boldsymbol{r}}_{i,t+1}$ $=(\hat{\boldsymbol{r}}_{i,t}N_{i,t}+\boldsymbol{r}_{a_{t},t}\cdot\mathds{1}_{\{a_{t}=i\}})/N_{i,t+1},\vspace{-5pt}$
Preference-aware optimization. To adapt the policy to the estimated preference, one might consider following the "optimism in the face of uncertainty" principle [17] by constructing confidence intervals on both preference and reward estimates, and incorporating them into the aggregation function $\Phi$ for optimization.
However, in this case, we claim that despite the deviation from the true expectation, a confidence term for the preference estimate $\hat{\boldsymbol{c}}_{t}$ is unnecessary The fundamental reason is that the preference estimation does not involve sequential action decision-making component, where the preference feedback ${\boldsymbol{c}}_{t}$ observed is independent of the chosen action $a_{t}$. Consequently, a bonus term for exploration is not required. In contrast, for reward estimation, the action $a_{t}$ determined by $\hat{\boldsymbol{r}}_{t}$ will also influence the future estimate $\hat{\boldsymbol{r}}_{t+1}$, making a confidence term on $\hat{\boldsymbol{r}}_{t}$ is necessary to encourage exploration.
Algorithm 1 Preference UCB with Unknown Preference (PRUCB-UP)
Initialize: $\alpha$, $N_{i,1}\!\leftarrow\!0$, $\boldsymbol{\hat{r}}_{i,1}\!\leftarrow\![0]^{D},\forall i\!\in\![K]$; $\boldsymbol{\hat{c}}_{1}\!\leftarrow\![0]^{D}$.
for $t=1,\cdots,T$  do
     Draw arm $a_{t}$ by (4),
     Observe reward $\boldsymbol{r}_{a_{t},t}$ and user preference $\boldsymbol{c}_{t}$.
     Update phase:
     Update preference estimate $\hat{\boldsymbol{c}}_{t+1}$ by (2); update $N_{i,t+1}$ and reward estimate $\hat{\boldsymbol{r}}_{i,t+1},\forall i\in[K]$ by (3).
end for
Building upon this, the arm selection policy is designed as:
$a_{t}=\operatorname*{arg\,max}_{i\in[K]}\Phi(\boldsymbol{\hat{c}}_{t},\hat{\boldsymbol{r}}_{i,t}+\rho_{i,t}^{\alpha}\boldsymbol{e}).\vspace{-2pt}$ (4)
where $\Phi$ is the inner product aggregation function, $\rho_{i,t}^{\alpha}=\sqrt{\log(t/\alpha)/\max\{1,N_{i,t}\}}$ is the standard Hoeffding bonus term. We characterize the regret of PRUCB-UP below.
Theorem 2.
Assume the preference follows unknown distribution with the value being revealed after each arm pull. Let $\eta_{i}=\boldsymbol{\overline{c}}^{T}\Delta_{i}$, $\Delta_{i}=\mu_{a^{*}}-\mu_{i}\in\mathbb{R}^{D}$, Algorithm 1 has
$R(T)=O\Bigg{(}\sum_{i\neq a^{*}}\Big{(}\underbrace{{\color[rgb]{0,1,1}\frac{\delta^{2}\log T}{\eta_{i}}+D\pi^{2}\alpha^{2}\eta_{i}}}_{R^{r}_{T}}+\underbrace{{\color[rgb]{1,.5,0}\frac{D^{2}\|\Delta_{i}\|_{2}^{2}\delta^{2}}{\eta_{i}}}}_{R^{c}_{T}}\Big{)}\Bigg{)},$
Remark 5.1.
Theorem 2 shows that with preference feedback, PRUCB-UP achieves a regret of $O(K\delta\log T+K\delta D^{2})$, demonstrating near-optimal performance. Notably, the regret caused by additional preference estimation error is bounded by a constant related to objective dimension $D$ and preference $\ell_{1}$-norm bound $\delta$. This implies that the impact of preference estimation error on the regret is small.
PRUCB-UP also admits an instance-independent regret bound. Due to space limit, we defer it to Appendix C.2. Additional, we show that the preference known case can be solved by PRUCB-UP as a special case, achieving a regret of $O(K\delta\log T)$, please see Appendix C.3 for details.
To prove Theorem 2, the main difficulty lies in decoupling and capturing the effect of the joint error from both reward estimation and preference estimation on the final regret. To address this, we introduce a tunable parameter $\epsilon$ to quantify the accuracy of preference estimation $\boldsymbol{\hat{c}}_{t}$, and decompose suboptimal actions into two disjoint sets: (1) suboptimal pulls under sufficiently precise preference estimation and (2) suboptimal pulls under imprecise preference estimation. For set (1), we demonstrate that it can be transferred to a preference-known instance with a diminished overall-reward gap to the optimal arm. For set (2), we transfer the original set with joint error to a preference estimation deviation event using Lemma 9 (Appendix C.1.2), making it more tractable. Please refer to Appendix C.1 for the full proof.
6 The Case with Hidden Preference

### v2

@@ 1 Introduction
Multi-objective multi-armed bandit (MO-MAB) is an important extension of standard MAB (Drugan and Nowe 2013). In MO-MAB problems each arm is associated with a $D$-dimensional reward vector. In this environment, objectives could conflict, leading to arms that are optimal in one dimension, but suboptimal in others. A natural solution is utilizing Pareto ordering to compare arms based on their rewards (Drugan and Nowe 2013). Specifically, for any arm $i\in[K]$, if its expected reward $\boldsymbol{\mu}_{i}$ is non-dominated by that of any other arms, arm $i$ is deemed to be Pareto optimal. The set containing all Pareto optimal arms is denoted as Pareto front $\mathcal{O}^{*}$. Formally, $\mathcal{O}^{*}=\{i\mid\boldsymbol{\mu}_{j}\not\succ\boldsymbol{\mu}_{i},\forall j\in[K]\setminus i\}$, where $\boldsymbol{u}\succ\boldsymbol{v}$ holds if and only if $\boldsymbol{u}(d)>\boldsymbol{v}(d),\forall d\in[D]$. The performance is then evaluated by Pareto regret, which measures the cumulative minimum distance between the learner’s obtained rewards and rewards of arms within $\mathcal{O}^{*}$ (Drugan and Nowe 2013). However, achieving low Pareto regret alone overlooks that users ultimately seek options aligned with their individual preferences. As the example depicted in Fig. 1, given multiple Pareto optimal restaurants, one user may give a higher preference to quality, while another user may give a higher preference to affordibility. This means that user preferences need to be accounted for in the MO-MAB problem set up in order to choose the right solution on the Pareto front $\mathcal{O}^{*}$. This is the focus of this paper.
@@
Numerous MO-MAB studies have been conducted but most of them achieve Pareto optimality via an arm selection policy that is uniform across all users, which we refer to as a global policy. One representative line of research focuses on efficiently estimating the entire Pareto front $\mathcal{O}^{*}$, and the action is randomly chosen on the estimated Pareto front (Drugan and Nowe 2013; Turgay, Oner, and Tekin 2018; Lu et al. 2019; Drugan 2018; Balef and Maghsudi 2023). Another line of research transforms the $D$-dimensional reward into a scalar using a scalarization function, which targets a specific Pareto optimal arm solution without the costly estimation of entire Pareto front Drugan and Nowe (2013); Busa-Fekete et al. (2017); Mehrotra, Xue, and Lalmas (2020); Xu and Klabjan (2023). These studies construct the scalarization function in a user-agnostic manner, causing the target arm solution to remain the same across different users (see Appendix A for a more detailed related work discussion). However, simply achieving Pareto optimality using a global policy may not yield favorable outcomes, since, as mentioned earlier, users often have diverse preferences across different objectives. Consider Fig. 1(a), where two users with different preferences interact with a conversational recommender to choose a restaurant based on multi-dimensional rewards (e.g., price, taste, service). Clearly, restaurants A, B, and C are Pareto optimal, as none of their rewards are dominated by others. Previous research using a global policy would either randomly recommend a restaurant from A, B, or C, or select one based on a fixed global criterion to achieve Pareto optimality. However, while recommending a restaurant like B might lead to positive feedback from user-1, it is likely to result in a low reward rating from user-2, who prefers an economical meal, since restaurant B is expensive. In contrast, Fig. 1(b) illustrates that when the system accurately captures user preferences (e.g., user-1 prefers a tasty meal, while user-2 prefers a cheap meal), it can select options more likely to receive positive reward ratings from both users. Therefore, we argue that optimizing MO-MAB should be customized based on the user preferences rather than solely aiming for Pareto optimality with a global policy.
@@
To fill this gap, we introduce a formulation of MO-MAB problem, where each user is associated with a $D$-dimensional preference vector, with each element representing the user’s preference for the corresponding objective. Formally, in each round $t$, user incurs a stochastic preference $\boldsymbol{c}_{t}\!\in\!\mathbb{R}^{D}$. The player selects an arm $a_{t}$ and observes a stochastic reward $\boldsymbol{r}_{a_{t},t}\!\in\!\mathbb{R}^{D}$. We define the scalar overall-reward as the inner product of arm reward $\boldsymbol{r}_{a_{t},t}$ and user preference $\boldsymbol{c}_{t}$. The learner’s goal is to maximize the overall-reward accrued over a given time horizon. We term this problem as Preference-Aware MO-MAB (PAMO-MAB).
@@
While interactive user modeling and customized optimization cross multiple objectives present promising experimental results in some areas including recommendation (Xie et al. 2021), ranking (Wanigasekara et al. 2019), and more (Reymond et al. 2024), there are no theoretical studies on MO-MAB customization under explicit user preferences. Particularly, two open questions remain: (1) how to develop provably efficient algorithms for customized optimization under different preference structure (e.g., unknown preference, hidden preference)? (2) how does the additional user preferences impact the overall performance?
@@
  • We make the first effort to address the open questions above. Motivated by real applications, we consider PAMO-MAB under two preference structures: unknown preference with feedback, and unknown preference without feedback (hidden preference), with tailored algorithms that are proven to achieve near-optimal regret in each case. These approaches is built on a designed general algorithmic backbone that introduces two key components: preference estimation and preference-aware optimization, to enable effective learning and decision-making under preference-centric customization. The expressions of our results are in an explicit form that capture a clear dependency on preference. To the best of our knowledge, this is the first work that explicitly showcases the fundamental impact of user preference in the regret optimization of MO-MAB problems.
@@
  • For the general hidden preference case, we propose a novel near-optimal algorithm PRUCB-HP that addresses the unique challenges of hidden PAMO-MAB with two key designs: (1) A weighted least squares-based hidden preference learner, with weights set as the inverse squared $\ell_{2}$-norm of reward observations, to resolve the random mapping issue caused by random preferences, and (2) A dual-exploration policy with novel bonus design to balance the trade-off between local exploration for identifying better reward arms and global exploration for refining preference learning. Additionally, we show that the unknown preference case can be viewed as a special instance of the hidden setting. A simplified variant, PRUCB-UP, naturally emerges under this setup, reusing the same backbone with reduced uncertainty and simplified estimation, while still achieving near-optimal regret cross all users.
  • Extensive experiments consistently validate the effectiveness of our algorithms in estimating preferences and rewards online, as well as in optimizing the overall reward.
@@ 2 Problem Formulation
We consider MO-MAB with $N$ users, $K$ arms and $D$ objectives. At each round $t\in[T]$, each user $n\in[N]$ is presented with an arm set $\mathcal{A}_{t}^{n}\subseteq[K]$, which may differ across users and time. The learner chooses an arm $a_{t}^{n}$ for user $n$ and observes a stochastic $D$-dimensional reward vector $\boldsymbol{r}_{a_{t}^{n},t}\in\mathcal{R}\subseteq\mathbb{R}^{D}$, which we refer to as reward. For the reward, we make the following standard assumption:
Assumption 1 (Bounded stochastic reward).
@@
User preferences. At each round $t$, we consider each user $n$ to be associated with a stochastic $D$-dimensional preference vector $\boldsymbol{c}_{t}^{n}\in\mathcal{C}\subseteq\mathbb{R}^{D}$, indicating the user preferences across the $D$ objectives. We refer to this vector as preference for short. Specifically, we make the following assumptions:
Assumption 2 (Bounded stochastic preference).
For $t\in[T],d\in[D],n\in[N]$, each preference entry $\boldsymbol{c}_{t}^{n}(d)$ is independently drawn from a fixed distribution (either known or unknown) with mean $\boldsymbol{\overline{c}}^{n}(d)$ and variance $\sigma_{c^{n},d}^{2}$, satisfying $\boldsymbol{c}^{n}(d)\geq 0$, $\|\boldsymbol{c}_{t}^{n}\|_{1}\leq\delta$, $\sigma_{c^{n},d}^{2}\in[0,\sigma^{2}_{c}]$.
Assumption 3 (Independence).
For any $t\!\in\![T]$, $n\!\in\![N]$, $i\!\in\![K]$, $d_{1},d_{2}\!\in\![D]$, $\boldsymbol{r}_{i,t}(d_{1})$, $\boldsymbol{c}^{n}_{t}(d_{2})$ are independent.
Assumption 3 is common in real applications since $\boldsymbol{c}_{t}$ and $\boldsymbol{r}_{t}$ are inherently determined by independent factors: user characteristics and arm properties. For example, an individual user’s preferences do not influence a restaurant’s location, environment, pricing level, etc., and vice versa.
Preference-aware reward. We define an overall-reward as the inner product of arm’s reward and user’s preference, which models the user reward rating under their preferences. In each round $t$, for each user $n$, the overall-reward score $g_{a_{t}^{n},t}$ for the chosen arm $a_{t}^{n}$ is defined as:
$g_{a_{t}^{n},t}={\boldsymbol{c}_{t}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}.$ (1)
To evaluate the learner’s performance, we define regret as the cumulative gap in overall-reward between selecting the optimal arm for each user at each round and the actual learner’s policy across the entire user set:
$R(T)=\sum^{T}_{t=1}\sum^{N}_{n=1}\boldsymbol{\overline{c}^{n}}^{\top}(\boldsymbol{\mu}_{a^{n*}_{t}}-\boldsymbol{\mu}_{a_{t}^{n}}),$ (2)
$a^{n*}_{t}\!=\!\operatorname*{arg\,max}_{i\in\mathcal{A}_{t}^{n}}\boldsymbol{\overline{c}^{n}}^{\top}\boldsymbol{\mu}_{i}$ is the optimal arm for user $n$ at round $t$. The goal is to minimize the regret $R(T)$. We term this problem as Preference-Aware MO-MAB (PAMO-MAB).
@@ 3 A Lower Bound
In the following, we develop a lower bound (Proposition 1) on the defined regret for PAMO-MAB. Such a lower bound will quantify how difficult it is to control regret without preference-adaptive policies under PAMO-MAB. Firstly, we present a definition characterizing a class of MO-MAB algorithms that are ”preference-free”.
@@
Let $\boldsymbol{\mathrm{c}}^{\top}=\{\boldsymbol{c}_{1},\boldsymbol{c}_{2},...,\boldsymbol{c}_{t}\}\in\mathbb{R}^{D\times t}$ be the preference sequence up to $t$ episode with mean $\boldsymbol{\overline{c}}$. Let $\pi_{t}^{\mathcal{A}}$ be the policy of algorithm $\mathcal{A}$ at time $t$ for selecting arm $a_{t}$. Then $\mathcal{A}$ is defined as preference-free if its policy $\pi_{t}^{\mathcal{A}}$ is independent of $\boldsymbol{\mathrm{c}}^{\top}$ and $\boldsymbol{\overline{c}}$, i.e., $\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i\mid\boldsymbol{\mathrm{c}}^{\top},\boldsymbol{\overline{c}})=\mathbb{P}_{\pi_{t}^{\mathcal{A}}}(a_{t}=i)$ for all arms $i\in[K]$ and all episodes $t\in[T]$.
To our knowledge, most existing algorithms in theoretical MO-MAB studies (Drugan and Nowe 2013; Busa-Fekete et al. 2017; Xu and Klabjan 2023; Hüyük and Tekin 2021; Cheng et al. 2024) fall within the class of preference-free algorithms, which employ a global policy for arm selection, while neglecting users’ preferences.
@@
Assume an MO-MAB environment contains multiple objective-conflicting arms, i.e., $|\mathcal{O}^{*}|\geq 2$, where $\mathcal{O}^{*}$ is the Pareto Optimal front. Then, for any preference-free algorithm, there exists a subset of users with distinct preferences such that the regret $R(T)=\Omega(T)$.
@@
Proposition 1 shows that for PAMO-MAB problem with $|\mathcal{O}^{*}|\geq 2$, sub-linear regret is unachievable for preference-free algorithms. This is because, for any arm $i\in\mathcal{O}^{*}$ that is optimal in one user preference subset $\mathcal{C}^{+}$, there exists another user preference subset $\mathcal{C}^{-}$ where arm $i$ becomes suboptimal, while preference-free algorithms cannot adapt their policies to varying preference across the entire space $\mathcal{C}$. Please see Appendix C for the detailed proof. We therefore ask the following question: Can we design preference-adaptive algorithms that achieve sub-linear regret for PAMO-MAB? The answer is yes. In the following, we analyze PAMO-MAB under two scenarios: hidden preference and preference feedback provided, demonstrating that with preference adaptation, sub-linear regret can indeed be achieved.
4 General Case with Hidden Preference